OpenStudy (anonymous):

I chose graphing..? is that the easiest way? Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 5x + 3y = 8 y = −2x + 1

OpenStudy (anonymous):

...

OpenStudy (anonymous):

I personally prefer to isolate the y on both equations then set them as equal to each other to solve for x. Then plug x into eithe requation to get teh y value. Put the x and y value together and you have your solution. In my opinion, the method is really preference. Do which ever you can do easiest. b. 5x+3y=8 3y=8-5x y=(8-5x)/3 (8-5x)/3=-2x+1 8-5x=3(-2x+1)=-6x+3 -5x+6x=3-8 x=-5 y=-2x+1 y=-2(-5)+1 y=10+1 y=11 (-5,11)

OpenStudy (anonymous):

Of course, if you have a graphing calculator, graphing method is probably faster.

OpenStudy (anonymous):

what method are you using?

OpenStudy (anonymous):

I'm not sure what it's called to be honest. My method is a version of subsitution that I learned. find the y of one equation by isolating the y value and subsituting it in for the y of another.

OpenStudy (anonymous):

It's really the same method except the original method does not require you to move y to one side on both equations. Just plug in y from the bottom equation tha you gave into teh top equation in place of y.

OpenStudy (anonymous):

Using the original method of subsitution: 5x+3y=8 5x+3(-2x+1)=8 5x-6x+3=8 -x=5 x=-5 y=-2x+1 y=-2(-5)+1 y=10+1 y=11 (-5,11)